What is 3 to Radical 2 + 5 to Radical 6
2 to Radical 3 + 2. This is in the form of a fraction. So (3_/2) + (5_/6
The number 2 has the square root sign next to it. (2_/3) + 2= Not very good at drawing the Radical symbol but it is a 3/2+
5/6. With Radical sign next to the number 6
Square root or {Radical sign next to the 2} 2/3 Radical sign next to 3+ 2 with no radical sign

Answer by richard1234(7193)

You might want to rewrite your question, since I am bound to misinterpreting what you typed. Perhaps use the LaTeX syntax:

3\sqrt{2} ==>

\sqrt[3]{2} ==>

\frac{3}{\sqrt{2}} ==>